# ARMFlow: AutoRegressive MeanFlow for Online 3D Human Reaction Generation

### v1

@@ 5 Experiments
We evaluate our approach on two widely adopted benchmarks for text-conditioned human interaction synthesis: InterHuman [intergen] and InterX [interx]. InterHuman comprises 7,779 interaction sequences, while InterX provides 11,388 sequences, each annotated with 3 textual descriptions. Compared with the other noisy and small datasets like NTU-120 and CHI3D[ntu_rgbd120, chi3d], these two datasets are of better motion qualities and have solid and public metrics for evaluations for fair comparisons.
The InterHuman dataset is built upon the AMASS [amass] skeleton, which includes 22 joints with the root joint. Each joint is represented following the HumanML3D [a2m] convention as $\{\mathbf{p}_{g},\mathbf{v}_{g},\mathbf{r}_{6d}\}$, where $\mathbf{p}_{g}\in\mathbb{R}^{3}$ denotes global position, $\mathbf{v}_{g}\in\mathbb{R}^{3}$ global velocity, and $\mathbf{r}_{6d}\in\mathbb{R}^{6}$ the local 6D rotation. This results in a motion tensor $\mathbf{m}_{p}\in\mathbb{R}^{N\times 22\times 12}$, accompanied by a binary foot-contact indicator $fc\in\mathbb{R}^{2}$.
In contrast, InterX adopts the SMPL-X [smplx] skeleton, which has 55 articulated joints covering the body, hands, and facial regions, along with the root orientation. Each joint rotation and root orientation is encoded as $\mathbf{r}_{6d}$, while the root translation and rotation are represented by $\mathbf{t}_{r}$ and $\mathbf{r}_{r}$, respectively, forming $\mathbf{m}_{p}\in\mathbb{R}^{N\times 56\times 6}$.
@@
For reaction generation, we mainly focus on fidelity and semantic correspondence. To comprehensively evaluate the performance of our model, we adopt a suite of feature-space metrics by intergen. To assess the realism and fidelity of generated interactions, we compute the Fréchet Inception Distance (FID) and between the feature distributions of generated and ground-truth motions and their Diversity. To evaluate the semantic alignment between text prompts and generated motions, we use R-Precision and Multimodal Distance (MMDist), which measure the consistency between text input and generated motions. To assess generative quality beyond accuracy, we report Multimodality, quantifying the model’s ability to produce multiple plausible motions for the same text prompts.
@@
InterGen [intergen] $9.183^{\pm.174}$ $0.325^{\pm.004}$ $0.467^{\pm.004}$ $0.546^{\pm.005}$ $3.859^{\pm.001}$ $7.305^{\pm.047}$ $1.270^{\pm.023}$
in2IN [in2in] $7.913^{\pm.251}$ $0.362^{\pm.004}$ $0.504^{\pm.008}$ $0.589^{\pm.007}$ $3.832^{\pm.002}$ $7.709^{\pm.040}$ $1.165^{\pm.034}$
MLD [mld] $3.588^{\pm.076}$ $0.405^{\pm.005}$ $0.561^{\pm.007}$ $0.649^{\pm.006}$ $3.798^{\pm.002}$ $7.663^{\pm.040}$ $1.124^{\pm.048}$
ReGenNet [xu2024regennet] $\underline{2.930}^{\pm.052}$ ${0.362}^{\pm.005}$ ${0.513}^{\pm.005}$ ${0.605}^{\pm.004}$ $3.815^{\pm.001}$ ${7.582}^{\pm.064}$ $\textbf{1.737}^{\pm.020}$
InterMask [intermask] ${3.453}^{\pm.061}$ $\underline{0.451}^{\pm.007}$ $\underline{0.610}^{\pm.006}$ $\underline{0.701}^{\pm.005}$ $\underline{3.782}^{\pm.002}$ $\underline{7.710}^{\pm.046}$ $\underline{1.361}^{\pm.032}$
@@
InterGen [intergen] $0.238^{\pm.038}$ $0.352^{\pm.004}$ $0.542^{\pm.005}$ $0.643^{\pm.004}$ $4.212^{\pm.029}$ $8.773^{\pm.067}$ $1.552^{\pm.029}$
MLD [mld] $0.148^{\pm.020}$ $0.414^{\pm.004}$ $0.607^{\pm.006}$ $0.712^{\pm.004}$ $3.655^{\pm.020}$ $8.893^{\pm.053}$ $\underline{1.875}^{\pm.088}$
ReGenNet [xu2024regennet] $\underline{0.093}^{\pm.022}$ $0.407^{\pm.004}$ $0.589^{\pm.004}$ $0.705^{\pm.003}$ $3.762^{\pm.015}$ $8.841^{\pm.072}$ $\textbf{1.937}^{\pm.069}$
InterMask [intermask] $0.399^{\pm.013}$ $\underline{0.429}^{\pm.005}$ $\underline{0.622}^{\pm.005}$ $\underline{0.731}^{\pm.005}$ ${3.584}^{\pm.017}$ ${8.911}^{\pm.057}$ $0.859^{\pm.033}$
@@
We divide our evaluation into two parts. For offline generation, we compare our approach with several state-of-the-art methods, including In2IN, which first synthesizes actor–reactor motions with an MDM prior and subsequently refines them; ReGenNet, a diffusion model with a transformer-decoder backbone; InterMask, a discrete autoregressive model based on masking; and an extended variant of MLD that performs diffusion in a latent space. For online generation, we benchmark against InterFormer, a conventional transformer-based autoregressive architecture; the online configuration of ReGenNet; and CAMDM, which adopts a transformer encoder and conditions on a fixed-length context window. Notably, these online baselines are not explicitly designed for real-time synthesis. Although ReGenNet can apply causal masking, its decoder architecture only conditions on the actor’s previous frame and cannot exploit the reactor’s past states, which often leads to temporal discontinuities. CAMDM, on the other hand, relies on a short fixed window of 10 frames, limiting long-range contextual modeling and causing pronounced drift in extended sequences, particularly on InterHuman. Furthermore, these designs do not support generation from scratch, rendering inference at t=0t{=}0 infeasible. In contrast, our method operates at a finer 4-frame granularity for real-time synthesis, scales more favorably than CAMDM, and does not depend on any initial motion. To ensure a fair comparison, we provide CAMDM with additional support by supplying an initial motion sequence, which compensates for its inability to generate from scratch, while keeping its 10-frame context window as specified by the authors.
@@
For the CNN encoder, we adopt the same design across both datasets except for the input dimension. Specifically, we follow the T2M-GPT configuration with a hidden size of 256, two residual convolutional downsampling blocks, and three hidden layers per block. For the main model, we use a standard DiT architecture as the backbone for both ARMFlow (online) and ReMFlow (offline), consisting of 512 hidden dimensions, 7 transformer layers, and 8 attention heads with skip connections. The MLP-based velocity predictor in ARMFlow is implemented as a lightweight 5-layer network, and we keep the network configuration identical across datasets. For MeanFlow, we sample timesteps from a logit-normal distribution and set the probability of instantaneous velocity sampling ($r=t$) to 0.25. In ReMFlow, classifier-free guidance (CFG) strength is set to $\omega=1.8$ for InterHuman and $\omega=2.0$ for InterX. For ARMFlow, we use $\omega=1.8$ on InterHuman and $\omega=1.2$ on InterX. All models are trained with a batch size of 64. For the CNN-VAE, we train for 2000 epochs on InterHuman and 800 epochs on InterX. ReMFlow is trained for 800 epochs on InterHuman and 500 epochs on InterX, while ARMFlow converges faster, requiring 500 epochs and 300 epochs, respectively. All experiments use AdamW optimizer with a learning rate of $1\times 10^{-4}$. All training and inference are conducted on NVIDIA H100 GPUs in an HPC environment.
@@ 5.1 Quantative Results
We first focus on the more challenging online generation setting. Tab. 2 reports the comparison between our ARMFlow and existing online generation models, ReGenNet and CAMDM. On the long-sequence InterHuman dataset, both baselines fail to produce satisfactory results. The primary limitation of CAMDM lies in its short context window, which prevents it from encoding the entire historical information, leading to severe temporal drift in the generated motions. Although ReGenNet utilizes the full actor history, its design initializes the reactor input from pure Gaussian noise at every inference step, making it unable to adaptively refine the reactor’s representation based on its own history, thus resulting in temporally inconsistent actions. In contrast, ARMFlow significantly outperforms existing methods in both FID and semantic consistency. Compared with CAMDM (the second-best method in FID), ARMFlow achieves a 45% relative improvement; and compared with ReGenNet (the second-best in semantic consistency), our model attains a 24.2% relative gain in R-precision (top-1). Moreover, on the shorter-sequence InterX dataset, our approach also achieves state-of-the-art results. ReGenNet performs relatively well on this dataset, which we attribute to its short-term attention mechanism being more effective when motion sequences are shorter. It is worth noting that online generation does not necessarily lead to performance degradation—in our experiments, ARMFlow even yields a lower (better) FID in the online setting than in the offline one, while the offline version, benefiting from full actor-conditioned information, achieves slightly higher semantic consistency.
@@
For the offline generation task, we further evaluate the generality of our ReMFlow and compare it with the current SOTA approaches in Tab. 3. InterMask, which is based on discrete random masking, performs well in semantic-matching metrics such as R-precision and MM-Distance, but its FID is considerably worse. This degradation mainly stems from its shared VQ-VAE encoder, which over-compresses motion features and thereby harms reconstruction quality—one of the key reasons we avoid discrete representations. ReGenNet, operating directly on raw data without compression, avoids such information loss; however, in the long-sequence InterHuman dataset, its self-attention mechanism struggles to capture fine-grained semantics, leading to weaker semantic consistency despite a good FID score. MLD lies between these two extremes, further validating a general trend in motion generation: stronger compression improves semantic alignment but often sacrifices reconstruction fidelity.
Notably, InterMask achieves its high semantic consistency at the cost of efficiency—it requires at least 20 inference steps, resulting in a total inference time of 0.77 s. In contrast, ReMFlow reaches state-of-the-art performance with a single forward pass, substantially improving inference speed while maintaining generation quality. Similar observations hold on the InterX dataset, where the shorter motion sequences allow ReGenNet to achieve relatively good semantic alignment as well.
@@ 5.3 Ablation Study
Table 6: Ablation on the training strategies for online autoregressive diffusion. GTE stands for ground-truth encoding, Rollout is the strategy used in HumanX[ji2025humanx], and BSCE is our strategy.

### v2

@@ 5 Experiments
We evaluate our approach on two widely adopted benchmarks for text-conditioned human interaction synthesis: InterHuman [21] and InterX [43]. InterHuman comprises 7,779 interaction sequences, while InterX provides 11,388 sequences, each annotated with 3 textual descriptions. Compared with the other noisy and small datasets like NTU-120 and CHI3D[24, 7], these two datasets are of better motion qualities and have solid and public metrics for evaluations for fair comparisons.
The InterHuman dataset is built upon the AMASS [27] skeleton, which includes 22 joints with the root joint. Each joint is represented following the HumanML3D [12] convention as $\{\mathbf{p}_{g},\mathbf{v}_{g},\mathbf{r}_{6d}\}$, where $\mathbf{p}_{g}\in\mathbb{R}^{3}$ denotes global position, $\mathbf{v}_{g}\in\mathbb{R}^{3}$ global velocity, and $\mathbf{r}_{6d}\in\mathbb{R}^{6}$ the local 6D rotation. This results in a motion tensor $\mathbf{m}_{p}\in\mathbb{R}^{N\times 22\times 12}$, accompanied by a binary foot-contact indicator $fc\in\mathbb{R}^{2}$.
In contrast, InterX adopts the SMPL-X [28] skeleton, which has 55 articulated joints covering the body, hands, and facial regions, along with the root orientation. Each joint rotation and root orientation is encoded as $\mathbf{r}_{6d}$, while the root translation and rotation are represented by $\mathbf{t}_{r}$ and $\mathbf{r}_{r}$, respectively, forming $\mathbf{m}_{p}\in\mathbb{R}^{N\times 56\times 6}$.
@@
For reaction generation, we mainly focus on fidelity and semantic correspondence. To comprehensively evaluate the performance of our model, we adopt a suite of feature-space metrics by Liang et al. [21]. To assess the realism and fidelity of generated interactions, we compute the Fréchet Inception Distance (FID) and between the feature distributions of generated and ground-truth motions and their Diversity. To evaluate the semantic alignment between text prompts and generated motions, we use R-Precision and Multimodal Distance (MMDist), which measure the consistency between text input and generated motions. To assess generative quality beyond accuracy, we report Multimodality, quantifying the model’s ability to produce multiple plausible motions for the same text prompts.
@@
InterGen [21] $9.183^{\pm.174}$ $0.325^{\pm.004}$ $0.467^{\pm.004}$ $0.546^{\pm.005}$ $3.859^{\pm.001}$ $7.305^{\pm.047}$ $1.270^{\pm.023}$
in2IN [33] $7.913^{\pm.251}$ $0.362^{\pm.004}$ $0.504^{\pm.008}$ $0.589^{\pm.007}$ $3.832^{\pm.002}$ $7.709^{\pm.040}$ $1.165^{\pm.034}$
MLD [4] $3.588^{\pm.076}$ $0.405^{\pm.005}$ $0.561^{\pm.007}$ $0.649^{\pm.006}$ $3.798^{\pm.002}$ $7.663^{\pm.040}$ $1.124^{\pm.048}$
ReGenNet [44] $\underline{2.930}^{\pm.052}$ ${0.362}^{\pm.005}$ ${0.513}^{\pm.005}$ ${0.605}^{\pm.004}$ $3.815^{\pm.001}$ ${7.582}^{\pm.064}$ $\textbf{1.737}^{\pm.020}$
InterMask [16] ${3.453}^{\pm.061}$ $\underline{0.451}^{\pm.007}$ $\underline{0.610}^{\pm.006}$ $\underline{0.701}^{\pm.005}$ $\underline{3.782}^{\pm.002}$ $\underline{7.710}^{\pm.046}$ $\underline{1.361}^{\pm.032}$
@@
InterGen [21] $0.238^{\pm.038}$ $0.352^{\pm.004}$ $0.542^{\pm.005}$ $0.643^{\pm.004}$ $4.212^{\pm.029}$ $8.773^{\pm.067}$ $1.552^{\pm.029}$
MLD [4] $0.148^{\pm.020}$ $0.414^{\pm.004}$ $0.607^{\pm.006}$ $0.712^{\pm.004}$ $3.655^{\pm.020}$ $8.893^{\pm.053}$ $\underline{1.875}^{\pm.088}$
ReGenNet [44] $\underline{0.093}^{\pm.022}$ $0.407^{\pm.004}$ $0.589^{\pm.004}$ $0.705^{\pm.003}$ $3.762^{\pm.015}$ $8.841^{\pm.072}$ $\textbf{1.937}^{\pm.069}$
InterMask [16] $0.399^{\pm.013}$ $\underline{0.429}^{\pm.005}$ $\underline{0.622}^{\pm.005}$ $\underline{0.731}^{\pm.005}$ ${3.584}^{\pm.017}$ ${8.911}^{\pm.057}$ $0.859^{\pm.033}$
@@
We divide our evaluation into two parts. For offline generation, we compare our approach with several state-of-the-art methods, including In2IN, which first synthesizes actor–reactor motions with an MDM prior and subsequently refines them; ReGenNet, a diffusion model with a transformer-decoder backbone; InterMask, a discrete autoregressive model based on masking; and an extended variant of MLD that performs diffusion in a latent space. For online generation, we benchmark against InterFormer, a conventional transformer-based autoregressive architecture; the online configuration of ReGenNet; CAMDM, which adopts a transformer encoder and conditions on a fixed-length context window; and R2R, an autoregressive diffusion diagram. Notably, these online baselines are not explicitly designed for real-time synthesis. Although ReGenNet can apply causal masking, its decoder architecture only conditions on the actor’s previous frame and cannot exploit the reactor’s past states, which often leads to temporal discontinuities. CAMDM, on the other hand, relies on a short fixed window of 10 frames, limiting long-range contextual modeling and causing pronounced drift in extended sequences, particularly on InterHuman. Furthermore, these designs do not support generation from scratch, rendering inference at t=0t{=}0 infeasible. In contrast, our method operates at a finer 4-frame granularity for real-time synthesis, scales more favorably than CAMDM, and does not depend on any initial motion. To ensure a fair comparison, we provide CAMDM with additional support by supplying an initial motion sequence, which compensates for its inability to generate from scratch, while keeping its 10-frame context window as specified by the authors.
@@
For the CNN encoder, we use the same architecture for both datasets except for the input dimension. Following T2M-GPT, the encoder has a hidden size of 256, two residual convolutional downsampling blocks, and three hidden layers per block. The main model adopts a standard DiT backbone for both ARMFlow (online) and ReMFlow (offline), with 512 hidden dimensions, 7 transformer layers, and 8 attention heads with skip connections. The MLP-based velocity predictor in ARMFlow is a lightweight 5-layer network, with identical settings across datasets. For MeanFlow, timesteps are sampled from a logit-normal distribution, with the probability of instantaneous velocity sampling ($r=t$) set to 0.25. In ReMFlow, classifier-free guidance (CFG) is set to $\omega=1.8$ for InterHuman and $\omega=2.0$ for InterX. For ARMFlow, we use $\omega=1.8$ on InterHuman and $\omega=1.2$ on InterX. All models are trained with a batch size of 64. The CNN-VAE is trained for 2000 epochs on InterHuman and 800 on InterX. ReMFlow is trained for 800 and 500 epochs, while ARMFlow converges faster with 500 and 300 epochs, respectively. All experiments use the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and are conducted on NVIDIA H100 GPUs in an HPC environment.
@@ 5.1 Quantative Results
We first consider the more challenging online generation setting. Tab. 2 compares our ARMFlow with existing online models, ReGenNet and CAMDM. On the long-sequence InterHuman dataset, both baselines fail to achieve satisfactory results. CAMDM is mainly limited by its short context window, which prevents encoding the full history and leads to severe temporal drift. Although ReGenNet leverages the full actor history, it initializes the reactor input from Gaussian noise at each inference step, preventing adaptive refinement of the reactor representation based on its own history and resulting in temporally inconsistent actions. In contrast, ARMFlow significantly outperforms existing methods in both FID and semantic consistency. Compared with R2R (the second-best in FID), ARMFlow achieves a 28% relative improvement. Moreover, on the shorter-sequence InterX dataset, our method also achieves state-of-the-art performance. ReGenNet performs relatively well on this dataset, likely because its short-term attention mechanism is more effective for shorter sequences. Notably, online generation does not necessarily degrade performance—in our experiments, ARMFlow even achieves a lower (better) FID in the online setting than in the offline one, while the offline version, benefiting from full actor-conditioned information, yields slightly higher semantic consistency.
@@
For the offline generation task, we evaluate the generality of ReMFlow and compare it with current SOTA methods in Tab. 3. InterMask, based on discrete random masking, performs well on semantic metrics such as R-precision and MM-Distance but yields a significantly worse FID. This mainly results from its shared VQ-VAE encoder, which over-compresses motion features and degrades reconstruction quality—one reason we avoid discrete representations. ReGenNet operates directly on raw data without compression and thus avoids this loss; however, on the long-sequence InterHuman dataset its self-attention struggles to capture fine-grained semantics, leading to weaker semantic consistency despite a strong FID. MLD lies between these extremes, reflecting a common trend in motion generation: stronger compression improves semantic alignment but often reduces reconstruction fidelity. InterMask achieves high semantic consistency at the cost of efficiency, requiring at least 20 inference steps (0.77 s). In contrast, ReMFlow attains state-of-the-art performance with a single forward pass, greatly improving inference speed while maintaining generation quality. Similar trends appear on the InterX dataset, where shorter sequences allow ReGenNet to achieve relatively strong semantic alignment.
@@ 5.3 Ablation Study
Table 6: Ablation on the training strategies for online autoregressive diffusion. GTE stands for ground-truth encoding, Rollout is the strategy used in HumanX[17], and BSCE is our strategy.
@@ 6 Conclusion
Acknowledgements
This research was supported by the Australian Research Council (ARC) Discovery Project DP240101926. We gratefully acknowledge this support.